### v18

@@ 1. Introduction
The ideas here are meant to give a general procedure for calculating an optimal sufficient value for the number of labels required in a standard labeling problem. The techniques apply to vertex labeling and edge labeling problems. Usually, the calculation will give a rough estimate of the optimal sufficient value, rather than an exact value. However, the calculation should bound the optimal value from above in most applications.
There are several steps to proving the main result, which is that given three sufficient conditions, it is possible to calculate upper bounds on the number of labels required in the general labeling problem. The first sufficient condition is that there is a uniform map that carries the fibers to one another transitively. This condition requires that the order on the permutations of the fibers is In most cases, this last equality will follow from the phrasing of the problem, but in the case that the labeling is on the vertices of the graph, and duplications are counted on induced labels on the edges, and the fibers of induction are the edges of the graph under the labeling scheme, the equality for requires that the graph under the labeling scheme is acyclic. Then, in addition, there is a basis step to an inductive argument. Usually, this step is verified by applying Markovβs Inequality to an expectation that is calculated for each of the individual fibers, before performing the identification allowed by the existence of the uniform map. Once the basis step is verified, apply induction.
The induction implies that there is a simple test for the verification of a fixed upper bound on number of required labels: The required step here is verification of the reference set hypothesis; show that the reciprocal of the sample set for the number of labelings is less than times the probability that there is a labeling without collisions. The probability, based on an expectation parameterized by the projected required strength of the labeling, is calculated using integration given by the induction from the basis step of the induction. There is a correlation between the size of an expected value, and the existence of a labeling without collisions, in the framework of each individual labeling problem.
@@
Step 3. In the reformulation of the problem, this step holds with and Take and use Markovβs Inequality. See that if then the basis step is satisfied. To check the reference set hypothesis, use . Count the number of labelings with replacement of both graphs. Then is about and is so this checks asymptotically. (It is required that be chosen slightly smaller than ) Try to use the edges as the fibers of induction, and get a function that is hyper-exponential in some base, and the form of is only
@@
For the -body problems, use the edges of the smaller graph as the fibers of induction; it is usually necessary to use the comparison orbit and maintain that the smaller graph in the placement is acyclic. In order to perform the basis step with use a comparison orbit of cardinality with an acyclic graph, to reconcile the basis step with an edge by edge resolution of a graph whose edges form the only fibers in the induction. Then, in all these cases, it is usually necessary to take the penalty for vertex label duplication. Try to extend the Erdos-Sos result to a -factor, using the edges as fibers, then the reference set hypothesis will not check unless The necessary restriction on in this example excludes any new results, and gives standard results within accepted parameters. Similarly, use the vertices as fibers and get results for cycles, and show gives and gives Try to extend the Sumner Conjecture to cycles. The phrasing of the problem as a vertex fiber problem will not permit a check of the reference set hypothesis.
@@
Summarize some of the techniques that are available for edge-disjoint placement problems. First, to pack for generally check the degree sequence condition, use a covering space with a multigraph, and possibly use a transformation on the random variable. (In this case, the graphs have to be acyclic.) Then, to pack and there are several different formulations of the problem, most of which rely on the -body allowance in some way, and some of which do not require the degree sequence condition as an initial hypothesis. First, take the edges as the fibers of induction. In this case, take the penalty for vertex label duplication and use the comparison orbit Do not check the degree sequence requirement. To apply this case, the graph has to be acyclic. Next, use the idea from the example on the Bollobas-Eldridge Conjecture.
In this case, check the degree sequence requirement explicity, but add to the expectation to account for the vertex label duplication because of the phrasing of the problem. Third, use the vertices as fibers, with the added penalty for vertex label duplication, with similar conditions to the edge fiber case, so that neither graph has to be acyclic, but formulate in terms of and take Additionally, in this case, the penalty for vertex duplication is based on instead of the order of the graph Here, it is not necessary to check the degree sequence condition separately. Fourth, take the vertices of one graph as the fibers of induction, and skip checking the degree sequence requirement. In this case, set the comparison orbit equal to the product of the cardinalities of the vertex sets of the individual graphs:
@@ 2. Degree sequences and edge-disjoint placements
The isometric embedding of the comparison orbit permits that all the fibers can be taken identically under the product topology. In there are useful identities for using binomial expansion, and when the isomorphic embedding is applied, the various identities form the basis step of the argument. Additionally, there is the very important use of the local topology in place of the standard graph model. Establish a set of maps from the multigraph or the local topology of the multigraph to the complete graph or the local topology of the complete graph respectively. Instead of using the model of at an edge, it is possible to look at the model at each vertex.
If there is not such that then every set of maps on takes a neighborhood of into for some Then for the product topology derived to be symmetric under so that every fiber can be taken identically (so that ), it is necessary to take That is, there is that realizes a packing so that the degree sequence is majorized by a lambda-fold complete graph with lambda taken greater than 1. Use this model for the set of maps, then use all the results in the setting of with the results will not apply to the simple complete graph, and will not give necessarily edge-disjoint placements. An example will make the idea clearer. Then the following theorem applies to trees whose degree sequences pack in The approach is to label the edges and vertices of a multigraph and then use the comparison orbit in the multigraph labeling problem as it isometrically translates into a separate model. Then use the basis step from the second model, where all the calculations for beta carry isometrically. The calculations in the second model cover the first model, so that when there are no duplications across all the fibers in the final step of the induction, this statement holds in the first model. So the final step of the induction in the second model implies there are no duplications across the fibers of the second model.
The conclusions only hold for trees, because the comparison orbit in the first model is only bounded by for acyclic graphs. There are two modifications for the problem. First, use a different model of the fibers than the complete graph. That is, use the model from the local topology of the complete graph. Second, because a separate model is used apart from the usual setting, there is an allowed transformation of the variable that makes the counting different.
@@
The standard form of gives a factor of in front of the term, however, the fibers double cover the space, so there is an additional factor of two in the expression for Then before the identification, and as the identification changes the random variable, so that after the identification. See that the non-negativity of the variance is not used here. The transformation of the variable is such that where is the product of two identical, independent random variables, so that before the identification. Apply the identification, and rescale the value of the random variable Then in these cases, checking the degree sequence condition will bound the magnitude of The degree sequence condition will be the main constraint in applying the technique. There are cases where the reference set hypothesis does not check; in this case, the reference set hypothesis checks. In this setting, take where is the number of fibers in because the local topology of applies instead of the model in the complete graph. The cardinality of the comparison orbit is given by Count, for the value times the cardinality of the coordinates in the edge labels on Then the first coordinate is fixed by the edge incident to the oriented edge on an out-directed copy of for all except when is incident from , the source of the out-directed copy of . The choice of the first label is arbitrary with uniform distribution. The probability and the event is the event that there are no edge-overlaps in
@@
The degree-sequence packing result is not going to show packing results in general. In the case that all the sizes of the elements of the packing are identical, with a threshold of where is a linear constant that is asympotically independent of the basis step checks, but the graphs may not be acyclic and the method is not going to apply. (This usage of is independent of the other usage of In some cases, the number of fibers in an edge-disjoint placement is not going to match the number of fibers in the covering space of the local topology. Trim down to an exact cover in these cases: The generating function with the binomial distibution as a check on the basis step does not apply, as in the fashion of the example. Again, when the model applies as the target for the fibers of the problem, the probability that there is a duplication on an element of the union of the targets (fibers) covers the probability that there is a duplication in the labeling. Generally speaking, take for the results to apply to simple graphs.
@@ 3. The dynamic reference set method
Begin an attempted proof that graceful labelings exist for all classes of trees. The attempted proof is based on the reference set method, but the proof here is limited to trees. The modification of the reference set method is called the dynamic reference set method. The proof is based on total probability, instead of inclusion-exclusion. The fibers of induction are still identified in the dynamic version of the method. However, the probability space expands as the induction continues, in this version of the method, so that the calculations are different. The domain of the fibers increases, under induction, and the probability space expands. In the primary case here, the labeling is on the vertices of the graph and the random variable counts duplicated labels induced on the edges of As with the regular version of the reference set method, there must be a resolution of the graph so that the fibers resolve one by one. Meanwhile, try counting the duplications of an edge labeling at the vertices of the graph: then the vertices have to account for the duplications at both endpoints of each edge. That is, the counting for is different if the vertices are used as fibers instead of edge fibers. The results that follow here use the edges as the fibers of induction, and, as a consequence, apply only to trees and forests. Start by considering Use to express the mean of a gamma random variable in this case.
The random variable that solves the related differential equation will be a simple gamma random variable in one parameter. Usually, use for that parameter, but in this case use values from the set The overall method is analogous to the reference set method in this case. Identify the fibers in the same way, but use total probability conditions on the counting for the fibers, instead of using inclusion-exclusion. The induction in this case is similar to the standard case, and works on larger and larger numbers of fibers at each step. That is, the argument is similar, but not identical to the standard argument. Most of the proof that
is identical. Induct on the probabilities using total probability instead of the usual inclusion-exclusion argument. The argument here is still quite general, but does not follow the same rules as the argument in the setting of inclusion-exclusion. In order to be complete, review the differential equation that generates the counting functions. In this case, see that
Usually the following applies:
In the case here, there are not the added complications with expressing the solutions, because of the different formulation of the problem. Rearranging the differential equation gives that where is the probability that the number of duplicates is less than or equal to on the sum of variables. This differential equation shows that the probability density function that gives the number of free labelings is approximated closely by the gamma distribution in one variable. In this case, take
To motivate this definition, check the expression for . The counting is complicated by the requirement that for all in order to solve the differential equation. If this equation is not satisfied, find and then reset for all values This generates consistent solutions as long as for all
If that is not the case, then set larger than such that for all and set for all Use these substitutions in every case of the application of the dynamic reference set method. The induction verifies that as the domain of the vertices of expands, from one to the random variable has
as the relevant inequality. For the main example, work with a labeling problem with the edges as fibers, and domain of on the vertices.
The constant is over some resolution of That is, count the cardinality of the reference set equal to to verify the reference set hypothesis. Multiply the reciprocal of by one in order to verify the reference set hypothesis.
In some cases, the assumption that for some realization of breaks down when is not a tree. Consider a case where the labeling duplications occur on the edges, and the edges are used as the fibers of induction. Instead of adding a pendant vertex, add a vertex with two chords. The inductive orbits of the fibers differ. For example, if there is a vertex with two chords after resolving three fibers, the function for in the validation of the reference set hypothesis is
In the tree case, there are five fibers, which are all interchangeable, including the choice of the fiber in the initial stage of the induction, and all the secondary fibers have interchangeable orders of resolution. In the case where one of the fibers necessarily has chords, then there are only four inductive orbits of fibers, and one of the inductive orbits has two inductive sub-orbits. In the acyclic (tree) case the counting is clear. In the case where the graph under the labeling scheme has cycles, the counting is more convoluted, but it is clear the two cases (acyclic as opposed to cyclic) do not necessarily have identical outcomes.
When the fibers of induction are vertices, there appears to be an analogous obstruction to the inductive step in the technique, but it does not appear to be identical. For instance, in a problem where the fibers of induction are vertices, but the label duplications are on edges, the dynamic version does not appear to apply, unless the duplications at each edge are counted at each of their endpoints. This complicates the basis step of the induction.
When the fibers of induction are vertices, but the label duplications are on vertices, the dynamic version appears to apply.
For instance, in the case of Brooksβ Theorem, the basis step would normally check, but in the context of the dynamic version of the reference set method, it is not the case that So in the case of Brooksβ Theorem, the dynamic version does not appear to apply.
(Example. The Graceful Tree Conjecture) So then, validating the inductive step and in turn the reference set hypothesis will then be markedly different in these cases where we add vertices with chords instead of pendant vertices.
@@
Step 2. The fibers of induction have to be edges, otherwise the count for the expectation will double for the number of endpoints of edges, in addition to the natural doubling of expectation that generally takes place. (The penalty portion of the expectation does double.) Calculate and then use to calculate the rest of the
Here, under the graceful labeling model, set and for where indexes the number of edges labeled, then set
for Fix for the sake of the argument, and see that there is a summand in the expectation for exceeding the range and that it is calculated correctly. Take
Then for all
Reset
as described. Assume that vertex labels are used only once, so there is no cycling.
Step 3. Define the random variables so that they count the number of duplicate labels on as the domain of expands. The expectations calculated are greater than or equal to the corresponding expectations for these random variables. To verify the basis case of the induction, verify that That is, show because is the cardinality of the solution space of and is the value of the probability distribution function for evaluated at 1. Use Markovβs Inequality and counting to get that
is sufficient to verify the basis case. The critical point is in the neighborhood of Then check the reference set hypothesis. Check Inequality (2) to verify the reference set hypothesis. Then the expression is roughly equivalent to
and it appears that this expression verifies the reference set hypothesis for the example is a tree of sufficiently large order.
It may be helpful to consider the cases for unions of cycles, and -graphs that are not acyclic.
To continue, first notice that there are different types of units on the random variable especially when the fibers are vertices that measure edge label duplications per vertex. In these cases, it seems to be that if there are edges per vertex, then is the probability that the duplications at independent edges, all rooted at the same vertex, are all zero. Then the edge label duplications will be zero more or less independently, depending on the problem.
@@ 4. Points of Errata
In all the cases where edge label duplications are used and the fibers are vertices, there is generally a fixed map from edges to vertices that establishes the vertex-root of an edge. In these cases, the condition does not necessarily hold unless the problem is properly phrased. Usually, the factorial in the denominator of the left-hand side of the check of the reference set hypothesis is accompanied by an identical factorial in the denominator of the right-hand side of the check of the reference set hypothesis, and these identical terms cancel, so that there is an abbreviated check of the reference set hypothesis. In order to do this, the orbits of the fibers are required to accumulate multiplicatively to form the factorial factor.
Try packing graphs or, equivalently, consider a labeling problem where the duplicates are between graphs instead of among graphs. The cardinality of the comparison orbit in is the total number of categories for fiber duplication, in these cases. That is, the comparison orbit is the number of categories which are considered for counting duplicate pairs. (In some cases, this is one-sided because the relation of a comparison orbit is not transitive: If and it is not necessarily the case that It is the case that the cardinalities of the comparison orbits do not have to be identical either.)
Resolve two cycles as in the case of -labelings. Then the factorial count for the order of the permutations of the fibers is not necessarily . The concern may be that the number of inductive orbits on the resolution of the graph is path-dependent, and that no resolution gives The probability is established by accounting for the number of edges mapped to the fiber under the map, as long as the fibers are vertices. If the fibers are edges, then the inductive orbits at a fiber of the resolution may fall into inductive sub-orbits, and this can cause some of the hypotheses of the reference set method to fail. The maximum factorial count for two disjoint cycles should be In this case, do not just apply this value to the denominator of the term in the reference set hypothesis check. Because the value is not exactly the reference set method does not apply at all. For this reason, it appears that the -labeling and -labeling results apply only to trees and forests (for all orders and sizes identically). It appears that in vertex labeling problems, when duplicates are counted on edges, and the edges are used as the fibers, the comparison orbit is not well-defined for cyclic graphs. That is, it appears that in these problems, the comparison orbits are only well-defined when the elements of the labeling problem are limited to acyclic graphs.
That is, there is an obstruction for vertex labeling problems, with vertices as fibers, where is taken over edges. In this case, the fibers can induce duplication units so that the ratio of fibers to edges cannot necessarily be taken in the case of unicyclic graphs. These restrictions do not apply to the Vizing Theorem or the Total Labeling Conjecture. In neither of these two cases do the labels on vertices induce labels on edges, though in the one case the labels on vertices may affect the comparison orbit on the edges.
The Gyarfas-Lehel result does not appear to apply to graphs generally, because, even though there is a separate model for the results, and the labeling problem is different, the comparison orbit is only defined as described for acyclic graphs, the label on the endpoint of each edge has possible values.
@@
Then the argument on the Tree Packing Conjecture has several complications that are not present in many of the other examples. Because there is not a clear example of a constructive technique that verifies the conclusions here regarding the Tree Packing Conjecture, these results should be viewed very critically. The proof included here on the Gyarfas-Lehel Conjecture assumes the local topology without fixing both the endpoints of the fibers. This technique of using the local topology is different from the other examples. Using a uniform map and a covering from the local topology of the complete graph, the events in the two distributions can be mapped The conclusions given about the conjecture should follow because there is an isomorphism between the two local topologies, and because the cardinality of the comparison orbit is carried isometrically from model to model. Apply the conclusion on the one model in the other model. The second model would normally have a static comparison orbit and the basis step is easier to do in the second model. Analogously, the cardinality of the comparison orbit is in the case of Bollobas-Eldridge, Erdos-Sos and Sumnerβs Conjecture if vertex-replacement is not allowed. However, in these three cases replacement is allowed and so the comparison orbit is Then the Erdos-Sos and Sumner Conjecture results apply to trees, but not necessarily other graphs. (In the case of Erdos-Sos, check the results that hold in the case ; the analysis suggests that there are some results.) The set-up for Gyarfas-Lehel gives a closed form for for a model of the problem labeling the edges of a multigraph (and some vertices) where is some packing of the graphs in a multigraph whose degree sequence is majorized by
In most cases, the reference set method gives a set of three initial conditions that are sufficient to demonstrate that the amount exceeds the threshold
@@
when is any non-negative integer. Consider a sequence of functions that converges point-wise to with discrete masses at Using weight the discrete points 1 at each discrete point in and on every point of the support
Take the limit of the sequence of functions as . Then the point-wise limit of the sequence is infinitely differentiable on the full support in the following sense: The limit of the sequence of functions is a cumulative density function under normalization, and converges point-wise to . Furthermore, the functions can be approximated by dirac-delta functions to get the differentiability of the cumulative distributions in the sense that (under proper realization) so that for all and upon induction,
Take as the normalizer of . That is, let be taken equal to the integral of over the full support with respect to Then if
it follows that, for instance,
as long as all the conditions from Step 1 and Step 2 of the reference set method are satisfied. The idea is that on some small support

### v19

@@ 1. Introduction
The ideas here are meant to give a general procedure for calculating an optimal sufficient value for the number of labels required in a standard labeling problem. The techniques apply to vertex labeling and edge labeling problems. Usually, the calculation will give a rough estimate of the optimal sufficient value, rather than an exact value. However, the calculation should bound the optimal value from above in most applications. There are several steps to proving the main result, which is that given three sufficient conditions, it is possible to calculate upper bounds on the number of labels required in the general labeling problem. The first sufficient condition is that there is a uniform map that carries the fibers to one another transitively. Then, in addition, there is a basis step to an inductive argument. Usually, this step is verified by applying Markovβs Inequality to an expectation that is calculated for each of the individual fibers, before performing the identification allowed by the existence of the uniform map. Once the basis step is verified, apply induction. The induction implies that there is a simple test for the verification of a fixed upper bound on the number of required labels. The required step here is verification of the reference set hypothesis; show that the reciprocal of the sample set for the number of labelings is less than the probability that there is a labeling without collisions. The probability, based on an expectation parameterized by the projected required strength of the labeling, is calculated using integration given by the induction from the basis step of the induction. There is a correlation between the size of an expected value, and the existence of a labeling without collisions, in the framework of each individual labeling problem.
@@
Step 3. In the reformulation of the problem, this step holds with and Take and use Markovβs Inequality. See that if then the basis step is satisfied. To check the reference set hypothesis, use . Count the number of labelings with replacement of both graphs. Then the value is about and is Try to use the edges as the fibers of induction, and get a function that is hyper-exponential in some base, whereas is only when replacement is allowed.
@@
For the -body problems, use the edges of the smaller graph as the fibers of induction, and usually maintain that the smaller graph in the placement is acyclic. In all these cases, it is usually necessary to take the penalty for vertex label duplication, and when the edges are the fibers of induction, the penalty doubles. To apply Markovβs Inequality to the basis case, compare some form of given by the individual fibers to the value The calculation for will generally vary from problem to problem, and is not always easy to establish. Try to extend the Erdos-Sos result to a -factor, using the edges as fibers. Then the reference set hypothesis will not check unless The necessary restriction on in this example excludes any new results, and gives standard results within accepted parameters. Similarly, use the vertices as fibers and get results for cycles. Show gives and gives Try to extend the Sumner Conjecture to cycles. The phrasing of the problem as a vertex fiber problem will not permit a check of the reference set hypothesis.
@@
Summarize some of the techniques that are available for edge-disjoint placement problems. First, to pack for generally check the degree sequence condition, use a covering space with a multigraph, and possibly use a transformation on the random variable. (In this case, the graphs have to be acyclic.) Then, to pack and there are several different formulations of the problem, most of which rely on the -body allowance in some way, and some of which do not require the degree sequence condition as an initial hypothesis. First, take the edges as the fibers of induction. In this case, take the penalty for vertex label duplication and use the cardinality of the comparison orbit Do not check the degree sequence requirement. To apply this case, the graph has to be acyclic. Next, use the idea from the example on the Bollobas-Eldridge Conjecture.
In this case, it does not appear necessary to check the degree sequence requirement explicity, but add to the expectation to account for the vertex label duplication because of the phrasing of the problem. Third, use the vertices as fibers, with the added penalty for vertex label duplication, with similar conditions to the edge fiber case, so that neither graph has to be acyclic, but formulate in terms of and take Additionally, in this case, the penalty for vertex duplication is based on instead of the order of the graph Here, it is not necessary to check the degree sequence condition separately. Fourth, take the vertices of one graph as the fibers of induction, and skip checking the degree sequence requirement. In this case, set the cardinality of the comparison orbit equal to the product of the cardinalities of the vertex sets of the individual graphs:
@@ 2. Degree sequences and edge-disjoint placements
The isometric embedding of the comparison orbit permits that the cardinality of the comparison orbit is well-defined and is calculated correctly. In there are useful identities for using binomial expansion, and when the isomorphic embedding is applied, the various identities form the basis step of the argument. Additionally, there is the very important use of the local topology in place of the standard graph model. Establish a set of maps from the multigraph or the local topology of the multigraph to the complete graph or the local topology of the complete graph respectively. Instead of using the model of at an edge, it is possible to look at the model at each vertex.
If there is not such that then every set of maps on takes a neighborhood of into for some That is, in some instances, every multigraph that realizes a packing is majorized by a lambda-fold complete graph with lambda taken greater than 1. Try using this model for the set of maps, so that all the results in the setting of have the results will not apply to the simple complete graph, and will not give necessarily edge-disjoint placements. If there is such that with then there is a valid covering space using the local topology of the complete graph. In this case, invoke induction using the cardinality of the comparison orbit, so that the reference set hypothesis applies. An example will make the idea clearer. The following theorem applies to trees whose degree sequences pack in The approach is to label the edges and vertices of a multigraph and then use the comparison orbit in the multigraph labeling problem as it isometrically translates into a separate model. Then use the basis step from the second model, where all the calculations for beta carry isometrically. The calculations in the second model cover the first model, so that when there are no duplications across all the fibers in the final step of the induction, this statement holds in the first model. That is, the final step of the induction in the second model implies there are no duplications across the fibers of the first model. The conclusions only hold for trees, because the comparison orbit in the first model is only bounded by for acyclic graphs. There are two modifications for the problem.
First, use a different model of the fibers than the complete graph. That is, use the model from the local topology of the complete graph. Second, because a separate model is used apart from the usual setting, there is an allowed transformation of the variable that makes the counting different.
@@
The standard form of gives a factor of in front of the term, however, the fibers double cover the space, so there is an additional factor of two in the expression for Then before the identification, and as the identification changes the random variable, so that after the identification. See that the non-negativity of the variance is used here. The transformation of the variable is such that where is the product of two copies of an identical random variable, so that before the identification. Apply the identification, and rescale the value of the random variable Then in these cases, checking the degree sequence condition will bound the magnitude of The degree sequence condition will be the main constraint in applying the technique. There are cases where the reference set hypothesis does not check; in this case, the reference set hypothesis checks. In this setting, take where is the number of fibers in because the local topology of applies instead of the model in the complete graph. The cardinality of the comparison orbit is given by Count, for the value times the cardinality of the coordinates in the edge labels on Then the first coordinate is fixed by the edge incident to the oriented edge on an out-directed copy of for all except when is incident from , the source of the out-directed copy of . The choice of the first label is arbitrary with uniform distribution. The probability and the event is the event that there are no edge-overlaps in
@@
The degree-sequence packing result is not going to show packing results in general. In the case that all the sizes of the elements of the packing are identical, with a threshold of where is a linear constant that is asympotically independent of the basis step checks, but the graphs may not be acyclic and the method is not going to apply. In some cases, the number of fibers in an edge-disjoint placement is not going to match the number of fibers in the covering space of the local topology. Trim down to an exact cover in these cases: The generating function with the binomial distibution as a check on the basis step does not apply, as in the fashion of the example. Again, when the model applies as the target for the fibers of the problem, the probability that there is a duplication on an element of the union of the targets (fibers) covers the probability that there is a duplication in the labeling. Generally speaking, take for the results to apply to simple graphs.
@@ 3. The dynamic reference set method
The ideas behind the reference set method can be used to formulate a proof technique that is more flexible in some contexts than the standard method. The modification of the reference set method is called the dynamic reference set method. Set
When figuring the expectation from do not divide by the number of fibers, or multiply by the cardinality of the comparison orbit. Simply fix From this starting point, the argument is similar, but not identical to the standard argument. In order to be complete, review the differential equation that generates the counting functions. In this case, see that
where Usually the following applies (with normally taken as ):
The three steps still apply, but Step 1 is more important, because the dynamic reference method sometimes permits the application of to satisfy the basis step. The contrast in the two forms of the primary differential equation is reflected in the types of applications.
The idea behind the dynamic reference set method is that if a system has a total expectation of then the system can be modeled by a set of random variables with In the dynamic reference set method, the base case can have a non-trivial solution space, and this consideration changes the counting. That is, the dynamic reference set method sometimes permits a simpler check for the basis case, and the case when has many helpful applications. In addition to the fact that the lower bound for is sometimes smaller in the dynamic version, the basis case only requires that one fiber satisfy the basis step. The standard method requires that all fibers satisfy the basis step. The main differential equation uses total probability, so the fibers are not identified in exactly the same way. Check that
in order to verify the reference set hypothesis. Particularly in the case considered here, it is required that the comparison orbit is well-defined, which will require in some cases that the graph under resolution is acyclic. For most of the problems where the fibers are vertices, the hypothesis on the comparison orbit is satisfied.
For instance, in the case of Brooksβ Theorem, the basis step of the dynamic reference set method does not check unless The inequality has to use the value taken On the other hand, when the value is sometimes too large to derive results for Brooksβ Theorem. Consider what happens when the value for approaches zero, and approaches some value equal to
(Example. The Graceful Tree Conjecture) Try to apply the dynamic reference set method to the popular conjecture about graceful labelings. The application here is limited to graphs that are acyclic.
@@
Step 2. The fibers of induction have to be edges, otherwise the count for the expectation will double for the number of endpoints of edges, in addition to the natural doubling of expectation that generally takes place. (The penalty portion of the expectation does double.) Calculate and then use to calculate the rest of the Here, under the graceful labeling model, the total penalty for vertex duplication is Set Measure
Fix for the sake of the argument. Use the expectation of the edge degree in the test for the basis step. The application of the expectation is permitted in the dynamic reference set method, because the labeling scheme is uniform to the labels, and the basis step has to hold for at least one of the fibers in the graph. Vary the value of in the interval The maximum value for is attained at the left end-point of the interval. When the induction is performed, set The number of fibers is set equal to The value for
Step 3. Evaluate Check the basis step first. That is, verify that
for sufficiently large. The inequality is closest when is maximized. Since the labels are uniform, take the two probabilities independently to get That is, the calculation for is independent of the labels chosen for the vertices that form the end-points of an arbitrary edge. The value can be taken The test for the reference set hypothesis is weaker than
which holds for sufficiently large, since and The value is sufficient to check the basis case, and the reference set hypothesis checks under this condition, as well.
Notice that, in general, the test for the basis step is
It may be helpful to consider the cases for unions of cycles, and -graphs that are not acyclic. To continue, first notice that there are different types of units on the random variable especially when the fibers are vertices that measure edge label duplications per vertex. In these cases, it seems to be that if there are edges per vertex, then is the probability that the duplications at independent edges, all rooted at the same vertex, are all zero. Then the edge label duplications will be zero more or less independently, depending on the problem.
@@ 4. Points of Errata
In some particular labeling problems, the comparison orbits are only well-defined when the elements of the labeling problem are limited to acyclic graphs. The Gyarfas-Lehel result does not appear to apply to graphs generally, because, even though there is a separate model for the results, and the labeling problem is different, the comparison orbit is only defined as described for acyclic graphs, the label on the endpoint of each edge has possible values. Then the argument on the Tree Packing Conjecture has several complications that are not present in many of the other examples. Because there is not a clear example of a constructive technique that verifies the conclusions here regarding the Tree Packing Conjecture, these results should be viewed very critically. The proof included here on the Gyarfas-Lehel Conjecture assumes the local topology without fixing both the endpoints of the fibers. This technique of using the local topology is different from the other examples. Using a uniform map and a covering from the local topology of the complete graph, the events in the two distributions can be mapped The conclusions given about the conjecture should follow because there is an isomorphism between the two local topologies, and because the cardinality of the comparison orbit is carried isometrically from model to model. Apply the conclusion on the one model in the other model. The standard model would normally have a static comparison orbit and the basis step is easier to do in the standard model. Analogously, the cardinality of the comparison orbit is in the case of Bollobas-Eldridge, Erdos-Sos and Sumnerβs Conjecture if vertex-replacement is not allowed. However, in these three cases replacement is allowed and so the comparison orbit is Then the Erdos-Sos and Sumner Conjecture results apply to trees, but not necessarily other graphs. (In the case of Erdos-Sos, check the results that hold in the case ; the analysis suggests that there are some results.) The set-up for Gyarfas-Lehel gives a closed form for for a model of the problem labeling the edges of a multigraph (and some vertices) where is some packing of the graphs in a multigraph whose degree sequence is majorized by The reference set method gives a set of initial conditions that are sufficient to show that the amount exceeds the threshold
@@
when is any non-negative integer. Consider a sequence of functions which are everywhere zero on except point masses at and the point mass where Weight the points 1, at each of the point masses, with on and zero elsewhere on the support. On the function Take the limit of the sequence of functions as . The point-wise limit of the sequence of functions is infinitely differentiable, in the sense that it is a cumulative density function under normalization, and the limit under is
That is, the functions can be replaced by dirac-delta functions to get the differentiability of a cumulative distribution in the sense that (under proper realization) so that for all and upon induction,
The distribution on for is not explicitly defined. The distribution on the will vary, and the only requirement that needs to be satisfied is that It follows that from the original differential equation on the Next, It follows that
as long as all the conditions from Step 1 and Step 2 of the reference set method are satisfied, and The idea is that on some small support
There appear to be several errors in [4]. For instance, in [4], it is suggested that
is sufficient to verify the reference set hypothesis in the dynamic case, and that whenever the comparison orbit is well-defined for the labeling problem, This ignores the possibility that and does not appear to be a useful way to think about the theory, because, for instance, in some cases and This last intended application still lends itself to interpretations where is some product of the cardinality of inductive orbits. That is, the product in the denominator of the right hand side of Inequality (2) is a -falling expression in some large value, where because the solution space of the base set is taken to be non-trivial in the dynamic reference set method. Then the ratio of on the left hand side of the proposed Inequality (3) is meant to be the ratio of the cardinality of the solution space of the base set with the product of cardinalities of the inductive orbits in [4]. However, see that the cardinality of the solution space of the base set is trivial in some cases, or much larger or smaller than so that this interpretation of the theory does not appear consistent. Rather, it appears that the Inequality (3) does not apply, but that the dynamic reference set method requires that the comparison orbit is well-defined, and that the product of the cardinality of the inductive orbits of a graph under resolution is
One of the advantages of the dynamic reference set method is that the base case only requires that one fiber pass the test for the basis case, as long as the labeling is uniform to the labels. That is, only one fiber has to pass the test for the basis case, as long as the explicit choice of labels applied to a fiber does not affect the probability that the label on the particular fiber has a duplication. Instead of using inclusion-exclusion, the dynamic reference set method uses total probability. Therefore, once the basis step is demonstrated for one fiber, the other fibers can be added sequentially. In order, subtract from the probability that the duplications over some set of fibers that contains the fibers in the induction has a label duplication, conditioned to the probability that the previously collected fibers do not register a duplication: Subtract The result is that the dynamic reference set method gives some resolution of a graph that has no label duplications.